OpenStudy (anonymous):

a triangle has the sides lenghts of 6,8,9... is it acute obtuse right or equiangular

OpenStudy (maheshmeghwal9):

wow interesting:D but i don't know how to do this:/

OpenStudy (anonymous):

haha i dont either

OpenStudy (anonymous):

the longer side will correspond to the larger angle. so the largest angle of this angle is correspond to the longest side which length is 9 for any triangle, if three sides are a,b,c and three angles are A,B,C we have a^2 = b^2 + c^2 - 2·b·c·cosA cosA = (c^2 + b^2 - a^2) / (2·b·c) u can use the length of three side to work out the largest angle and judge wheather it is

OpenStudy (anonymous):

9^2=6^2+8^2-2*6*8*cosA cosA=(9^2-6^2-8^2)/(2*6*8)=(-19)/96

OpenStudy (anonymous):

cosA is negative and A belongs to 0 to pi. so A is larger than 90 degrees. so, it is a obtuse triangle

OpenStudy (anonymous):

where did you learn all this stuff??? and thanks for help me

OpenStudy (anonymous):

well, it is just formula. i learn it from books

OpenStudy (anonymous):

your really smart

OpenStudy (anonymous):

:)